# Learning multi-modal representations by watching hundreds of surgical video lectures

### v3

@@ Abstract
Recent advancements in surgical computer vision applications have been driven by vision-only models, which do not explicitly integrate the rich semantics of language into their design. These methods rely on manually annotated surgical videos to predict a fixed set of object categories, limiting their generalizability to unseen surgical procedures and downstream tasks. In this work, we put forward the idea that the surgical video lectures available through open surgical e-learning platforms can provide effective vision and language supervisory signals for multi-modal representation learning without relying on manual annotations. We address the surgery-specific linguistic challenges present in surgical video lectures by employing multiple complementary automatic speech recognition systems to generate text transcriptions. We then present a novel method, SurgVLP - Surgical Vision Language Pre-training, for multi-modal representation learning. SurgVLP constructs a new contrastive learning objective to align video clip embeddings with the corresponding multiple text embeddings by bringing them together within a joint latent space. To effectively demonstrate the representational capability of the learned joint latent space, we introduce several vision-and-language surgical tasks and evaluate various vision-only tasks specific to surgery. Contrary to current surgical computer vision fully supervised approaches, we show that SurgVLP can be adapted to different surgical procedures and tasks without any specific fine-tuning. To the best of our knowledge, this is the first work to show the zero-shot adaptation to current vision-only surgical downstream tasks, such as surgical tool, phase, and triplet recognition, without any manual annotation. These results highlight the learned multi-modal representations’ transferability and versatility in surgical video analysis. The code is available at https://github.com/CAMMA-public/SurgVLP.
Keywords: Multi-modal representation learning; Surgical video lectures; Self-supervision; Vision-and-language.
@@ 1 Introduction
Next, we assess the robustness and adaptability of our approach when applied to unseen surgical datasets and tasks. Specifically, we examine its performance in traditional vision-only surgical tasks, including surgical tool, phase, and action triplet recognition [94, 67]. We evaluate our approach as a zero-shot transfer learning by processing the category labels (tool, phase, or action triplet) into textual form and classifying the video frames based on the similarity of visual and textual latent vectors. The results show that the general surgical concepts learned by our multi-modal joint representations from various surgical procedures can benefit a specific surgical procedure, such as laparoscopic cholecystectomy. To the best of our knowledge, this is the first work demonstrating the ability of a learned multi-modal representation to recognize surgical tools, phases, and action triplets without using any ground truth labels. While the results fall short of fully-supervised baselines, they are encouraging and could pave the way for more advanced methodologies to improve the results further. Finally, we conduct extensive ablation studies to shed light on the different components of our approach and their impacts on the results. The contributions of our work can be succinctly summarized in the following four key aspects:
@@ 2.2 Downstream tasks
The surgical computer vision community has embraced these advancements, exploring numerous studies on surgical tool and phase recognition [94, 23, 1, 89]. Following this, researchers proposed more refined spatiotemporal architectures to capture the dynamic and complex nature of surgical tasks [41, 20, 21, 42, 80]. For a more comprehensive understanding, a recent survey provides an overview of surgical phase recognition approaches [31]. For a more fine-grained surgical activity recognition, authors in [67] have introduced the challenging action triplet recognition task for laparoscopic cholecystectomy procedures. They propose a dataset to classify instrument, verb, and target components from surgical scenes. However, all previous methods are vision-only, fully supervised methods that rely heavily on human-annotated curated data.
@@ 3.1 Video clip-text pairs
We utilize the recent advancements in audio speech recognition (ASR) to transcribe lecture audio into textual sentences. In particular, we use two different ASR systems to generate complementary text views, namely the AWS Medical Transcribe ASR system [6] and the Whisper ASR system [75], as we observe that ASR systems perform differently on the specific surgical domain. In particular, AWS Medical Transcribe is used to understand medical terminology and surgery-specific terms, and Whisper is used to understand the overall sentence structure and common words.
@@ 5.5.1 Vision-and-language tasks
Fig. 5 shows the impact of novel text queries on the temporal activity grounding task. Specifically, we simulate the open-vocabulary scenario by decomposing the sentence into sub-sentences. For example, we construct a clear and concise text query by removing the sub-sentence “through umbilical incision” from the complete sentence “the bag be pull through the umbilical incision on the specimen be retrieve”, resulting in text query “the bag be pull and the specimen be retrieve”. We plot the visual-textual activation map by comparing the text embedding to all visual frame embeddings of the video. Also, as shown in Fig. 5 right, the combination of two sentences will generate the decentralized activation map that spreads the attention to different timestamps, indicating that the SurgVLP model is triggered to the unambiguous semantics, and removing the redundant text query can improve the temporal activity grounding performance. As shown in Fig. 6, SurgVLP trained with two text views generates a focused attention map, highlighting relevant frames and ignoring adjacent ones. This demonstrates that multiple text view supervision effectively connects the visual and the textual content. Compared to the AWS text view, the model pretrained with Whisper text view generates the attention map with focused areas when the text query is about the long-term video activity, such as “sequential application of bipolar cautery and cut with scissor”.
@@ 6.1 Discussion
This work shows that the proposed SurgVLP achieves superior zero-shot performance than the state-of-the-art methods [74] from the general computer vision field. The better performance is enabled due to the constructed large-scale surgical vision-language dataset and pretraining strategy with multiple text views. However, the zero-shot adaptation of SurgVLP is not supervised by any annotated data, leading to a suboptimal performance compared to the fully supervised works [94, 20]. A potential improvement toward real-world application is to adapt the learned multi-modal representations of the pretrained SurgVLP to the downstream tasks using less labeled data with fully-supervised finetuning. Specifically, the SurgVLP’s dual-branch architecture can encode domain-specific textual knowledge while capturing detailed visual patterns from the surgical scene [44]. Consequently, the feature extractor of fully supervised methods [94, 20] could be boosted by the complementary information from the textual side. Another future work line is to explore “cheaper” self-supervision signals within the visual and textual modalities. Typical work includes building the external knowledge base [90] and performing retrieval-augmented vision-language pretraining [102]. Witnessing the recent emergence of large language models [93] and its encoded clinical knowledge, exploring the usage of these language models by eliciting their knowledge can help to merge the domain gap. Furthermore, the current work overlooks the hierarchical structure inherent in surgical videos. To address this, hierarchical multi-modal pretraining can be incorporated to further improve the performance for the surgical downstream tasks requiring long temporal context for the prediction.”

### v5

@@ Abstract
Recent advancements in surgical computer vision applications have been driven by vision-only models, which do not explicitly integrate the rich semantics of language into their design. These methods rely on manually annotated surgical videos to predict a fixed set of object categories, limiting their generalizability to unseen surgical procedures and downstream tasks. In this work, we put forward the idea that the surgical video lectures available through open surgical e-learning platforms can provide effective vision and language supervisory signals for multi-modal representation learning without relying on manual annotations. We address the surgery-specific linguistic challenges present in surgical video lectures by employing multiple complementary automatic speech recognition systems to generate text transcriptions. We then present a novel method, SurgVLP - Surgical Vision Language Pre-training, for multi-modal representation learning. SurgVLP constructs a new contrastive learning objective to align video clip embeddings with the corresponding multiple text embeddings by bringing them together within a joint latent space. To effectively demonstrate the representational capability of the learned joint latent space, we introduce several vision-and-language surgical tasks and evaluate various vision-only tasks specific to surgery, e.g., surgical tool, phase, and triplet recognition. Extensive experiments across diverse surgical procedures and tasks demonstrate that the multi-modal representations learned by SurgVLP exhibit strong transferability and adaptability in surgical video analysis. Furthermore, our zero-shot evaluations highlight SurgVLP’s potential as a general-purpose foundation model for surgical workflow analysis, reducing the reliance on extensive manual annotations for downstream tasks, and facilitating adaptation methods such as few-shot learning to build a scalable and data-efficient solution for various downstream surgical applications. The code is available at https://github.com/CAMMA-public/SurgVLP.
keywords:
Multi-modal representation learning , Surgical video lectures, Self-supervision , Vision-and-language
@@ 1 Introduction
Next, we assess the robustness and adaptability of our approach when applied to unseen surgical datasets and tasks. Specifically, we examine its performance in traditional vision-only surgical tasks, including surgical tool, phase, and action triplet recognition [94, 67]. We evaluate our approach as a zero-shot transfer learning by processing the category labels (tool, phase, or action triplet) into textual form and classifying the video frames based on the similarity of visual and textual latent vectors. The results show that the general surgical concepts learned by our multi-modal joint representations from various surgical procedures can benefit a specific surgical procedure, such as laparoscopic cholecystectomy. To the best of our knowledge, this is the first work to demonstrate self-supervised multi-modal pretraining for recognizing surgical tools, phases, and action triplets without annotation. While our zero-shot performance trails fully supervised baselines, particularly in tasks requiring fine-grained anatomical reasoning, the results highlight SurgVLP’s potential to serve as a foundation backbone to reduce annotation costs for downstream tasks. Finally, we conduct extensive ablation studies to shed light on the different components of our approach and their impacts on the results. The contributions of our work can be succinctly summarized in the following four key aspects:
@@ 2.2 Downstream tasks
The surgical computer vision community has embraced these advancements, exploring numerous studies on surgical tool and phase recognition [94, 23, 1, 89]. Following this, researchers proposed more refined spatiotemporal architectures to capture the dynamic and complex nature of surgical tasks [41, 20, 21, 42, 80]
For a more comprehensive understanding, a recent survey provides an overview of surgical phase recognition approaches [31]. For a more fine-grained surgical activity recognition, authors in [67] have introduced the challenging action triplet recognition task for laparoscopic cholecystectomy procedures. They propose a dataset to classify instrument, verb, and target components from surgical scenes. However, all previous methods are vision-only, fully supervised methods that rely heavily on human-annotated curated data.
@@ 3.1 Video clip-text pairs
We utilize the recent advancements in audio speech recognition (ASR) to transcribe lecture audio into textual sentences. In particular, we use two different ASR systems to generate complementary text views, namely the AWS Medical Transcribe ASR system [6] and the Whisper ASR system [75], as we observe that ASR systems perform differently on the specific surgical domain. In particular, AWS Medical Transcribe is used to understand medical terminology and surgery-specific terms, and Whisper is used to understand the overall sentence structure and common words. As shown in Table. 1, we have included a direct comparison of text generated by AWS and Whisper to illustrate their respective strengths in ASR for surgical language understanding. Our observations confirm that AWS excels in recognizing specific surgical terms, making it particularly useful for capturing domain-specific vocabulary accurately. For example, in the first case, AWS correctly transcribes “trocar,” which is an important surgical instrument, while Whisper misinterprets it as “choker.” However, Whisper provides a more structured and coherent transcription, as seen in the second example, where it accurately reconstructs the full sentence with proper context, compared to AWS’s fragmented output. These differences highlight a key trade-off: AWS is more precise in recognizing surgical terms, whereas Whisper generates more fluent and contextually complete sentences. This distinction is crucial when designing multi-view models, as combining outputs from both ASR systems helps leverage AWS’s terminology accuracy and Whisper’s sentence coherence, leading to more effective language supervision for surgical video-language pretraining.
AWS Whisper
this be possible through the use of a trocar connect to a glove this is possible through the use of a choker connected to a glove a tip that has been described numerous times in the literature
a tube
the plication be perform start on the right side of the esophagus
on use any thigh limb
two oh stitch
a two-way fund application is performed starting on the right side of the esophagus and using ethybon 2 0 stitches
Table 1: Comparison of transcriptions generated by AWS and Whisper ASR systems.
@@ 3.2.1 Visual encoder
In this work, we adopt ResNet-50 with ImageNet initialization as our visual encoder, prioritizing spatial feature extraction in surgical videos. While Vision Transformers (ViTs) have demonstrated strong capabilities in modeling long-range dependencies, their effectiveness heavily relies on large-scale pretraining datasets, which are often unavailable in specialized domains like surgical vision-language tasks. Moreover, previous works in surgical workflow analysis predominantly utilize CNN-based architectures, making ResNet-50 a fair baseline for comparison. Additionally, ViTs exhibit higher initialization sensitivity, leading to less stable performance in our vision-language pretraining framework. Since temporal modeling is beyond the scope of this work, we focus on extracting strong spatial representations, where ResNet-50 proves to be a robust and efficient choice for surgical video understanding.
@@ 5.5.1 Vision-and-language tasks
Fig. 5 shows the impact of novel text queries on the temporal activity grounding task. Specifically, we simulate the open-vocabulary scenario by decomposing the sentence into sub-sentences. For example, we construct a clear and concise text query by removing the sub-sentence “through umbilical incision” from the complete sentence “the bag be pull through the umbilical incision on the specimen be retrieve”, resulting in text query “the bag be pull and the specimen be retrieve”. We plot the visual-textual activation map by comparing the text embedding to all visual frame embeddings of the video. Also, as shown in Fig. 5 right, the combination of two sentences will generate the decentralized activation map that spreads the attention to different timestamps, indicating that the SurgVLP model is triggered to the unambiguous semantics, and removing the redundant text query can improve the temporal activity grounding performance. As shown in Fig. 6, SurgVLP trained with two text views generates a focused attention map, highlighting relevant frames and ignoring adjacent ones. This demonstrates that multiple text view supervision effectively connects the visual and the textual content. Compared to the AWS text view, the model pretrained with Whisper text view generates the attention map with focused areas when the text query is about the long-term video activity, such as “sequential application of bipolar cautery and cut with scissor”. Fig. 6 also demonstrates that incorporating multiple text views helps generate textual activation maps with reduced noise, suggesting a more refined multi-modal embedding space that better aligns images with semantically similar language descriptions. However, the observed misalignment with ground truth highlights a key limitation: temporal grounding depends on accurate boundary detection, which our pretrained video-text pairs cannot fully provide. This discrepancy arises because spoken descriptions often precede corresponding visual actions, creating an inherent latency between audio transcriptions and visual content. As a result, the activation maps may exhibit shifts in alignment, reflecting this temporal gap.
@@ 6.1 Discussion
6.1.1 Future Work
This work shows that the proposed SurgVLP achieves superior zero-shot performance than the state-of-the-art methods [74] from the general computer vision field. The better performance is enabled due to the constructed large-scale surgical vision-language dataset and pretraining strategy with multiple text views. However, the zero-shot adaptation of SurgVLP is not supervised by any annotated data, leading to a suboptimal performance compared to the fully supervised works [94, 20]. A potential improvement toward real-world application is to adapt the learned multi-modal representations of the pretrained SurgVLP to the downstream tasks using less labeled data with fully-supervised finetuning. Specifically, the SurgVLP’s dual-branch architecture can encode domain-specific textual knowledge while capturing detailed visual patterns from the surgical scene [44]. Consequently, the feature extractor of fully supervised methods [94, 20] could be boosted by the complementary information from the textual side. Another future work line is to explore “cheaper” self-supervision signals within the visual and textual modalities. Typical work includes building the external knowledge base [90] and performing retrieval-augmented vision-language pretraining [102]. Witnessing the recent emergence of large language models [93] and its encoded clinical knowledge, exploring the usage of these language models by eliciting their knowledge can help to merge the domain gap. Furthermore, the current work overlooks the hierarchical structure inherent in surgical videos. To address this, hierarchical multi-modal pretraining can be incorporated to further improve the performance for the surgical downstream tasks requiring long temporal context for the prediction.
6.1.2 Limitation
While SurgVLP reduces annotation burdens, its zero-shot performance remains below fully supervised methods, particularly in complex tasks requiring fine-grained anatomical reasoning or handling intraoperative variability. Also, reliance on error-prone ASR systems introduces transcription inaccuracies in surgical terminology and temporal misalignments between narrated steps and visual content, magnified by domain shifts between lecture demonstrations and real-world surgical environments. In addition, pretraining data lacks diversity for rare or specialized surgeries, while conventional accuracy metrics inadequately reflect clinical risks, e.g., error impacts on patient safety. Finally, computational costs may hinder deployment in resource-limited settings. These limitations position SurgVLP as a pretraining strategy to augment, but not replace, fully-supervised learning, requiring targeted fine-tuning and domain-adaptation methods for clinical applicability.